# A unified framework for geometry-independent operator learning in cardiac electrophysiology simulations

### v1

@@ 3 Discussion
The increasing complexity of modern biomedical simulations, particularly in fields like computational cardiology [brown2025cardiac], has created a growing need for methods that can bridge the gap between high-fidelity mechanistic models and the demands of real-time clinical applications. The results demonstrate that our proposed framework can accurately and robustly learn the complex, nonlinear mapping from a patient’s anatomy and electrophysiological parameters to their full-field LAT map, achieving high fidelity to biophysical simulations in milliseconds. This represents an advancement over conventional approaches and a crucial step toward enabling in-procedure, real-time, personalised cardiac electrophysiology analysis.
@@ 3.1 Geometry-independent neural operator learning
Our approach demonstrates that a well-designed coordinate system is a powerful tool for bridging the gap between irregular, real-world data and the structured-domain requirements of neural networks, allowing a single neural operator to learn a geometry-independent mapping. By leveraging the UAC system, we established a canonical representation that enabled neural operators to robustly capture the complex, nonlinear physics of cardiac conduction [hao2023gnot, li2023transformer, shih2025transformers, ovadia2024vito]. A key finding is the model’s low MAE of $(5.15\pm 0.25)$ ms. This level of accuracy is significant in a clinical context, as it is well within the typical measurement noise and physiological variability observed in cardiac electroanatomic mapping systems [gaeta2021high] (typically on the order of 5–10 ms [narayan2024advanced]). This demonstrates that our model’s predictions are not only numerically accurate but also fall within a clinically acceptable margin, making the approach viable for non-invasive patient stratification, procedure planning, and guiding ablation therapies.
A central tenet of our work is the understanding that the success of a neural operator relies on the availability of a large, diverse, and high-quality training dataset. By developing a GPU-accelerated FEM solver, we were able to generate a large simulation dataset across diverse patient-specific left atrial anatomies, providing a robust foundation for learning generalizable principles of cardiac electrophysiology. This data generation pipeline is a critical component of our framework, as it bypasses the prohibitive computational bottleneck that has historically limited the application of data-driven methods in this domain.
Beyond supporting our framework, this large, curated dataset provides a foundation for future data-driven studies in cardiac electrophysiology. Its scale, diversity, and consistency enable predictive models to generalise across anatomies and protocols, a capability previously limited by data scarcity. For comparison, G-FuNK [loeffler2024graph], trained on 25 left atrial meshes from atrial fibrillation patients, concluded that additional geometries could improve performance. DIMON [yin2024scalable] used 1,006 left ventricular patient data, and each was simulated at seven pacing sites, totalling 7,042 simulations. In contrast, our simulations preserve these locations while also varying tissue conductivity, allowing neural operators to capture electrophysiological responses across diverse stimulation patterns and heterogeneous, patient-specific left atrial anatomies. Overall, our dataset spans 308,700 simulations, offering a far richer foundation for training and benchmarking state-of-the-art neural operators.
@@ 3.3 Toward domain-invariant generalisation
This finding establishes a generalizable strategy for achieving domain-invariant medical image analysis. Our UAC projection provides the necessary geometry-invariant foundation, while training on heterogeneous data acts as a regularization method in feature space, compelling the model to learn only the biophysically relevant structures common to both domains and emphasizing that true translational power derives from integrating diverse, multi-centre data rather than from isolated single-site optimisation.
@@ 3.4 Limitations and future directions
Despite its strong performance, the present study has limitations. The observation that removing fibre orientation features improved model performance was initially counterintuitive. Our analysis suggests two primary reasons for this finding. First, the current representation of fibre orientation may contain inherent noise or inaccuracies that interfere with the model’s ability to learn the true underlying electrophysiological dynamics. Second, the model appears to be leveraging the spatial and geometric priors provided by the coordinates and surface area features more effectively. These geometric inputs may implicitly contain sufficient information about the overall anatomical structure, making the explicit fibre orientation data redundant or even counterproductive in its current form.
@@
Beyond the feature representation, limitations also arise from the use of the UAC system. The UAC system is designed to map the closed, complex 3D left atrial surface onto a flat, 2D grid. To achieve this flattening, a continuous line must be cut on the 3D surface. Depending on the chosen coordinate orientation, this seam manifests as a central line on the 2D map. This seam creates a spatial discontinuity for the learning model, as points that are physically adjacent on the 3D surface are severed across this line in the 2D domain. Consequently, the neural operator struggles to accurately interpolate the LAT function across this boundary, resulting in a consistent band of large error centred precisely along this necessary central UAC seam.
The additional large localised errors, which are more distributed, are plausibly linked to areas of extreme anatomical complexity and steep electrophysiological gradients. These clusters primarily occur around the highly patient-specific structures like the PV junctions and LAA. The UAC projection cannot fully remove the residual geometric variability present in these complex regions, increasing the model’s uncertainty. Furthermore, activation wavefronts often exhibit the steepest LAT gradients in these areas. The model is challenged to precisely localise and reproduce these sharp functional variations across diverse anatomies, leading to the small, perceptible clusters of higher error in these critical physiological locations. These difficulties suggest that while our model excels at capturing global activation patterns, the fine-grained, localised details influenced by complex anatomical geometries remain a challenging frontier. Future work could explore incorporating a more fine-grained representation of these critical regions to further enhance local accuracy.
@@
Finally, the framework was trained exclusively on high-fidelity simulated datasets. Although these simulations spanned a wide range of anatomies and electrophysiological parameters, direct validation against clinical electroanatomic mapping data is essential to confirm robustness under real-world conditions. The current implementation focuses on the left atrium; extending the approach to whole-heart or biventricular modelling will require adapting the anatomical projection and scaling the data generation process. Future directions include combining hybrid simulation–clinical datasets, refining anatomical embeddings, and exploring adaptive spatial resolutions to capture fine-scale electrophysiological detail without compromising inference speed.
@@ 4 Methods
We developed a unified, scalable computational framework that overcomes the computational bottlenecks of conventional EP simulations (Fig. 3). This framework is designed to circumvent the computational bottleneck inherent in conventional EP simulations. It leverages a diverse, high-fidelity dataset generated by a custom high-performance FEM solver to enable the training of a neural operator that can predict LAT maps with accuracy comparable to ground truth, but within milliseconds.
@@ 4.1.1 Datasets description
4.1.2 UAC projection
We projected the spatial coordinates and fibre directions on the mesh, along with the resulting LAT map, to structured grids of fixed size by UAC. The UAC system [uac] defines a bijective transformation
$\Phi:\mathcal{M}\rightarrow[0,1]^{2},$
where $\mathcal{M}\subset\mathbb{R}^{3}$ denotes the atrial surface manifold, and $0,1]^{2}$ is a normalised two-dimensional coordinate space. The map $\Phi$ is constructed using anatomical landmarks, typically the pulmonary vein junctions and the mitral annulus, as anchor points to define physiologically meaningful coordinates. These coordinates correspond to a geodesic distance-based unfolding of the atrial surface, such that anatomical regions (e.g., anterior wall, posterior wall, roof) consistently map to aligned regions of the 2D domain across patients.
The standardised representation allows to map topological entities from different anatomy to corresponding regions. Let $f:\mathcal{M}\rightarrow\mathbb{R}$ be a scalar function defined on the atrial mesh (e.g., LAT values), and let $(\alpha_{i},\beta_{i})=\Phi(\mathbf{x}_{i})$ be the UAC coordinates of a mesh vertex $\mathbf{x}_{i}$. We then construct a function $\tilde{f}:[0,1]^{2}\rightarrow\mathbb{R}$ via spatial interpolation:
$\tilde{f}(\alpha,\beta)=\mathcal{I}\left(\left\{(\alpha_{i},\beta_{i},f_{i})\right\}_{i=1}^{N}\right),$
where $\mathcal{I}(\cdot)$ denotes a suitable interpolation operator.
In practice, we use a hybrid of linear and nearest-neighbour interpolation to preserve continuity in densely sampled regions while maintaining robustness in sparsely sampled areas. The 3D vertex coordinates $(x,y,z)$ for a mesh are projected directly to the 2D grid of shape $(3,50,50)$ where the first dimension (3) represents the channels corresponding to the $x$, $y$, and $z$ coordinates, and the latter two dimensions ($50\times 50$) define the resolution of the structured grid in the UAC system.
Fibre orientations, however, are originally defined per mesh element rather than per vertex. To obtain vertex-wise fibre directions suitable for UAC projection, we convert element-level vectors to nodal vectors using area-weighted averaging. For each vertex, we accumulate the fibre vectors from all incident triangles, scaling each vector by the area of its corresponding triangle so that larger elements contribute proportionally. The resulting vectors are normalised to unit length to preserve directional information. These vertex-level fibre directions are then mapped to the 2D grid using the same UAC interpolation procedure as these coordinates, resulting in a 2D grid of shape $(3,50,50)$.
@@ 4.2 GPU-based FEM EP simulation
To generate a large-scale, comprehensive dataset for training, we simulated LA electrical propagation using the widely adopted monodomain equation [geselowitz1983bidomain] (further implementation details are provided in Supplementary Section 1). This biophysical model captures the electrical conduction through anisotropic cardiac tissue, balancing capacitive and ionic currents with diffusive current flow, governed by the longitudinal ($\sigma_{l}$) and transverse ($\sigma_{t}$) conductivities.
The primary and clinically significant output extracted from these simulations is the Local Activation Time (LAT) map, which defines the time $t$ at which the transmembrane potential $V$ first crosses a predefined threshold ($V_{\text{th}}$, typically 0 mV) at any location $x$ in the cardiac domain:
@@
Each simulation was run for 600 ms to capture the full activation sequence, and the local activation time (LAT) map was recorded during runtime. Each simulation required approximately one minute on an NVIDIA A100 GPU, enabling the generation of 210,000 simulations for Cohort A and 98,700 simulations for Cohort B, yielding a combined total of 308,700 simulations. Executing these simulations across eight A100 GPUs reduced the total wall-clock time from an estimated 214 days (single GPU) to approximately 27 days.

### v2

@@ 3.1 Geometry-independent neural operator learning
Our approach demonstrates that a well-designed coordinate system is an efficient tool for bridging the gap between irregular, real-world data and the structured-domain requirements of neural networks, allowing a single neural operator to learn a geometry-independent mapping. By leveraging the universal coordinate system, we established a canonical representation that enabled neural operators to robustly capture the complex, nonlinear physics of cardiac conduction [hao2023gnot, li2023transformer, shih2025transformers, ovadia2024vito]. The resulting prediction accuracy, with mean absolute errors on the order of 5 ms, lies within the typical measurement uncertainty and physiological variability reported for clinical electroanatomic mapping systems [gaeta2021high] (typically on the order of 5–10 ms [narayan2024advanced]). This demonstrates that our model’s predictions are not only numerically accurate but also fall within a clinically acceptable margin.
A central premise of this work is that effective neural operator learning in cardiac electrophysiology depends on access to large, diverse, and high-quality training data. To this end, we developed a GPU-accelerated finite element electrophysiology solver that enables efficient generation of high-fidelity simulations across a wide range of patient-specific left atrial anatomies and electrophysiological conditions. This data generation pipeline addresses a longstanding computational bottleneck in cardiac modelling and provides a scalable foundation for learning generalisable electrophysiological structure.
Beyond supporting our framework, this large, curated dataset provides a foundation for future data-driven studies in cardiac electrophysiology. Its scale, diversity, and consistency enable predictive models to generalise across anatomies and protocols, a capability previously limited by data scarcity. For context, G-FuNK [loeffler2024graph] was trained on 25 left atrial geometries, while DIMON [yin2024scalable] leveraged 1,006 ventricular geometries simulated at multiple pacing sites. In contrast, our simulations preserve these locations while also varying tissue conductivity, allowing neural operators to capture electrophysiological responses across diverse stimulation patterns and heterogeneous, patient-specific left atrial anatomies. Overall, our dataset spans 308,700 simulations, offering a far richer foundation for training and benchmarking state-of-the-art neural operators.
The observation that removing fibre orientation features improved model performance was initially counterintuitive. Our analysis suggests two primary reasons for this finding. First, the current representation of fibre orientation may contain inherent noise or inaccuracies that interfere with the model’s ability to learn the true underlying electrophysiological dynamics. Second, the model appears to be leveraging the spatial and geometric priors provided by the coordinates and surface area features more effectively. These geometric inputs may implicitly contain sufficient information about the overall anatomical structure, making the explicit fibre orientation data redundant or even counterproductive in its current form.
@@ 3.3 Toward domain-invariant generalisation
This finding establishes a generalizable strategy for achieving domain-invariant medical image analysis. The universal coordinate projection provides the necessary geometry-invariant foundation, while training on heterogeneous data acts as a regularisation method in feature space, compelling the model to learn only the biophysically relevant structures common to both domains and emphasising that true translational power derives from integrating diverse, multi-centre data rather than from isolated single-site optimisation.
@@ 3.4 Limitations and future directions
Beyond feature representation, limitations also arise from the use of intrinsic anatomical coordinate systems. For the left atrium, the UAC system maps a closed, topologically complex three-dimensional surface onto a flat two-dimensional domain. This flattening necessarily requires introducing a cut line on the surface, which appears as a seam in the coordinate representation. As a result, points that are physically adjacent on the original anatomy become separated across this seam, creating an artificial spatial discontinuity for the learning model. Consistent with this construction, our results reveal a narrow, coherent band of elevated prediction error aligned with the UAC seam, indicating that the neural operator struggles to interpolate smoothly across this boundary.
In addition to seam-related effects, further localised errors are observed in regions of pronounced anatomical and electrophysiological complexity, particularly around the pulmonary vein junctions and the left atrial appendage. These regions exhibit substantial inter-patient variability that cannot be fully normalised by the UAC projection, leading to residual geometric heterogeneity in the coordinate domain. Moreover, activation wavefronts in these areas often display steep local gradients in activation time, increasing sensitivity to small spatial misalignments. Together, these factors likely contribute to the small but systematic clusters of elevated error observed in physiologically critical regions. These observations indicate that while the proposed framework captures global activation patterns and long-range structure effectively, resolving fine-scale, anatomy-driven variations remains challenging.
By contrast, ventricular geometries parameterised using the UVC do not exhibit a comparable global seam-induced discontinuity. Instead, elevated errors are primarily confined to regions introduced by geometric truncation, such as the basal valve plane and the interface with the right ventricle. This contrast highlights that coordinate-induced limitations are strongly dependent on underlying anatomical topology rather than intrinsic to the operator-learning framework itself, and that intrinsic coordinate representations may behave differently across cardiac chambers.
@@
Finally, the framework was trained exclusively on high-fidelity simulated datasets. Although these simulations span a wide range of anatomies and electrophysiological parameters, validation against clinical electroanatomic mapping data will be essential to assess robustness under real-world conditions. In addition, the current implementation focuses primarily on atrial and ventricular geometries in isolation; extending the approach to whole-heart or fully biventricular modelling will require adapting the anatomical parameterisation and scaling the data generation process. Future directions include integrating hybrid simulation–clinical datasets, refining anatomical representations in regions of high complexity, and exploring adaptive spatial resolution strategies to further improve local accuracy without compromising inference efficiency.
@@ 4 Methods
We describe a computational framework that combines high-fidelity electrophysiology simulations with neural operator learning to enable efficient prediction of LAT maps (Fig. 4).
@@ 4.1.1 Datasets description
4.1.2 Universal coordinates projection
Spatial coordinates, fibre directions, and LAT fields from patient-specific mesh geometries are projected onto structured grids of fixed resolution using universal anatomical coordinate systems. For the left atrium, UAC defines a bijective transformation
$\Phi_{\mathrm{LA}}:\mathcal{M}\rightarrow[0,1]^{2},$
where $\mathcal{M}\subset\mathbb{R}^{3}$ denotes the atrial surface manifold. The mapping is constructed from anatomical landmarks, including the pulmonary vein junctions and mitral annulus, yielding a geodesic distance-based unfolding that aligns corresponding anatomical regions across patients [uac]. An analogous intrinsic parameterisation is applied to ventricular geometries using the UVC system, which maps the three-dimensional ventricular volume to a structured coordinate domain.
Let $f:\mathcal{M}\rightarrow\mathbb{R}$ denote a scalar field on the mesh (e.g., LAT), and let $\boldsymbol{\xi}_{i}=\Phi(\mathbf{x}_{i})$ be the corresponding UAC or UVC coordinates of mesh point $\mathbf{x}_{i}$. A coordinate-domain representation $\tilde{f}$ is constructed via spatial interpolation,
$\tilde{f}(\boldsymbol{\xi})=\mathcal{I}\!\left(\{(\boldsymbol{\xi}_{i},f_{i})\}_{i=1}^{N}\right),$
where $\mathcal{I}(\cdot)$ combines linear interpolation in densely sampled regions with nearest-neighbour interpolation near anatomical boundaries.
For atrial meshes, three-dimensional vertex coordinates $(x,y,z)$ are mapped to a grid in the UAC domain. Fibre orientations, originally defined per element, are converted to vertex-wise vectors using area-weighted averaging and normalised to unit length before projection. An equivalent procedure is applied in the UVC domain for ventricular geometries. After neural operator inference in the coordinate domain, predictions can be mapped back to the original patient-specific geometry. For each mesh point $\mathbf{x}\in\mathcal{M}$ with coordinate $\boldsymbol{\xi}=\Phi(\mathbf{x})$, the predicted value is recovered by interpolating the learned field $\hat{f}(\boldsymbol{\xi})$.
@@ 4.2 GPU-based FEM EP simulation
To generate a large-scale, comprehensive dataset for training, we simulated LA electrical propagation using the widely adopted monodomain equation [geselowitz1983bidomain] (further implementation details are provided in Supplementary Section 1, with an example of simulated wavefront propagation shown in Supplementary Fig. 1 and benchmark verification results in Supplementary Fig. 2). This biophysical model captures the electrical conduction through anisotropic cardiac tissue, balancing capacitive and ionic currents with diffusive current flow, governed by the longitudinal ($\sigma_{l}$) and transverse ($\sigma_{t}$) conductivities.
The primary and clinically significant output extracted from these simulations is the LAT map, which defines the time $t$ at which the transmembrane potential $V$ first crosses a predefined threshold ($V_{\text{th}}$, typically 0 mV) at any location $x$ in the cardiac domain:
@@
Each simulation was run for 600 ms to capture the full activation sequence, and the LAT map was recorded during runtime. Each simulation required approximately one minute on an NVIDIA A100 GPU, enabling the generation of 210,000 simulations for Cohort A and 98,700 simulations for Cohort B, yielding a combined total of 308,700 simulations. Executing these simulations across eight A100 GPUs reduced the total wall-clock time from an estimated 214 days (single GPU) to approximately 27 days.
@@ 4.3.2 Sobolev Loss with Spatial Smoothness Regularisations
Author contributions
B.Z. developed the computational methodology, performed simulations, and drafted the manuscript. C.C. contributed to model development and data analysis. S.Q. assisted with numerical experiments and model development. M.B. contributed to software implementation and optimisation. A.W.C.L. prepared the data. C.R. contributed to manuscript editing and provided the simulation data. C.R., M.J.W.G. and L.H.G.A.H. provided experimental data. M.Q. assisted with analysing results. S.N. supervised the study, contributed to conceptualisation, and revised the manuscript. All authors reviewed the manuscript.
Competing interests
The authors declare no competing interests.